# Adaptive Spatial Goodness Encoding: Advancing and Scaling Forward-Forward Learning Without Backpropagation

### v1

@@ Abstract
The Forward-Forward (FF) algorithm offers a promising alternative to backpropagation (BP). Despite advancements in recent FF-based extensions, which have enhanced the original algorithm and adapted it to convolutional neural networks (CNNs), they often suffer from limited representational capacity and poor scalability to large-scale datasets, primarily due to exploding channel dimensionality. In this work, we propose adaptive spatial goodness encoding (ASGE), a new FF-based training framework tailored for CNNs. ASGE leverages feature maps to compute spatially-aware goodness representations at each layer, enabling layer-wise supervision. Crucially, this approach decouples classification complexity from channel dimensionality, thereby addressing the issue of channel explosion and achieving competitive performance compared to other BP-free methods. ASGE outperforms all other FF-based approaches across multiple benchmarks, delivering test accuracies of 99.65% on MNIST, 93.41% on FashionMNIST, 90.62% on CIFAR-10, and 65.42% on CIFAR-100. Moreover, we present the first successful application of FF-based training to ImageNet, with Top-1 and Top-5 accuracies of 26.21% and 47.49%. By entirely eliminating BP and significantly narrowing the performance gap with BP-trained models, the ASGE framework establishes a viable foundation toward scalable BP-free CNN training.
Indroduction
@@
Artificial Neural Networks (ANNs) have become a cornerstone of modern machine learning, largely owing to the effectiveness of the backpropagation (BP) algorithm (Rumelhart, Hinton, and Williams 1986). BP has driven the success of ANNs across a diverse range of applications (Goodfellow et al. 2014; He et al. 2015; LeCun, Bengio, and Hinton 2015; Kingma and Ba 2017; Vaswani et al. 2023) by enabling efficient training through gradient-based optimization. The algorithm proceeds in two phases: a forward pass computes the network’s predictions and their associated error relative to the ground truth, followed by a backward pass that propagates this error through the network. This yields gradients for each layer’s parameters (weights and biases), which are then used by optimization methods such as stochastic gradient descent (SGD) (Bottou 2010) to iteratively minimize the prediction error. However, BP suffers from inherent limitations. It requires symmetric weight matrices for the forward and backward passes to compute parameter gradients via the chain rule (Grossberg 1987). Additionally, it necessitates the storage of intermediate activations until parameter updates occur (Czarnecki et al. 2017). These constraints incur substantial computational and memory overhead, while simultaneously restricting the degree of parallelization, ultimately rendering BP inefficient and ill-suited for deployment on resource-constrained hardware.
@@
In response, numerous backpropagation-free (BP-free) strategies have emerged to address these limitations, including Target Propagation (TP) (Bengio 2014), Feedback Alignment (FA) (Lillicrap et al. 2016), Direct Feedback Alignment (DFA) (Nøkland 2016), Deep Random Target Projection (DRTP) (Frenkel, Lefebvre, and Bol 2021), and Recursive Local Representation Alignment (rLRA) (Ororbia et al. 2023), alongside other biologically inspired or local learning methods (Hebb 2002; Moraitis et al. 2022; Dellaferrera and Kreiman 2023). Although these approaches alleviate some of the BP limitations, they continue to face challenges in balancing the trade-offs between prediction accuracy, memory efficiency and computational cost.
@@
A recent promising alternative is the Forward-Forward algorithm (FF) (Hinton 2022), which employs a layer-wise training strategy through dual forward passes: one with positive samples and the other with negative samples. Each layer independently updates its parameters to maximize a “goodness” metric for positive inputs while minimizing it for negative ones. By eliminating the need for backward passes, FF significantly reduces memory overhead and facilitates pipelined parallelism. However, Figure 1 illustrates a critical limitation of FF when applied to convolutional neural networks (CNNs): prediction accuracy tends to deteriorate progressively with network depth. This indicates a weakening capacity for feature abstraction in deeper layers, thereby constraining the algorithm’s effectiveness in deep architectures and on complex tasks. Recent enhancements, such as Channel-wise Competitive (CwC) learning (Papachristodoulou et al. 2024) and DeeperForward (Sun et al. 2025), attempt to mitigate this issue. While partially effective, they introduce an undesirable coupling between channel dimensionality and classification complexity. Specifically, they require layer-wise channels $C_{l}$ to scale linearly with classes $N$, i.e., $C_{l}\in\Theta(N)$. This scaling becomes impractical for large-scale datasets like ImageNet (Russakovsky et al. 2015) due to the dramatic increase in channel dimensionality.
@@
BP (Rumelhart, Hinton, and Williams 1986) enables end-to-end optimization by propagating errors from the output layer back through the network. However, it suffers from several inherent limitations. First, it requires weight symmetry between the forward and backward passes (Grossberg 1987). Second, BP introduces the update locking problem (Czarnecki et al. 2017), wherein weight updates cannot be performed until the entire forward pass is completed. These constraints result in high memory overhead due to the need to store intermediate activations and hinder layer-wise parallelism, ultimately rendering BP inefficient for deployment on resource-constrained hardware.
@@
To overcome the limitations of backpropagation, numerous alternatives have been proposed that either avoid or approximate its core mechanism. Feedback Alignment (FA) (Lillicrap et al. 2016) eliminates the requirement for weight symmetry in the backward pass by using fixed, randomly initialized feedback weights. DFA (Nøkland 2016) extends this by propagating output layer errors to hidden layers, bypassing the backward pathway altogether. Another related method, rLRA (Ororbia et al. 2023) enables weight updates without computing global gradients by recursively aligning each layer’s output with a synthetic target derived from the layer above. PEPITA (Dellaferrera and Kreiman 2023) replaces the backward pass with a second forward pass, in which the input is modulated according to the output-layer errors from the first forward pass. Notably, these approaches still necessitate full forward passes to compute global errors or store intermediate activations prior to updating weights. To eliminate update locking, local learning rules—which allow layers to update independently without waiting for the entire forward pass—are desirable. DRTP (Frenkel, Lefebvre, and Bol 2021) achieves this by directly projecting true labels to each layer, generating layer-wise error signals that support local updates without waiting for the full forward pass. Another approach is Hebbian learning (Hebb 2002), a biologically inspired local learning rule that strengthens synaptic connections when presynaptic and postsynaptic neurons are activated simultaneously. Despite its biological plausibility, Hebbian learning does not support supervised learning or error correction. SoftHebb (Moraitis et al. 2022; Journé et al. 2023) extends Hebbian learning by incorporating soft winner-take-all dynamics and differentiable surrogate gradients, enabling the training of deeper networks while preserving the characteristics of local learning. Similarly, Predictive Coding (Rao and Ballard 1999; Millidge, Seth, and Buckley 2022) offers a local learning framework in which each layer predicts the activity of the layer below, using the resulting prediction error to locally update both prediction and representation neurons.
@@
The FF algorithm (Hinton 2022) introduces a fundamentally different training paradigm that entirely avoids the need for backpropagation. It employs a layer-wise learning mechanism in which each layer is independently supervised through two distinct forward passes—one with positive data and the other with negative data. Each layer computes a “goodness” score, defined as the sum of squared activations, and is trained using a binary cross-entropy loss to classify inputs as positive if the goodness exceeds a predefined threshold, and negative otherwise. Several extensions have been proposed to enhance the performance of FF. Contrastive Forward-Forward (Aghagolzadeh and Ezoji 2025) redefines goodness as the dot-product similarity between representations of different samples and introduces layer-wise supervised contrastive learning to improve the discriminative power of the network. SymBa (Lee and Song 2023) addresses the inherent asymmetry in binary cross-entropy loss by introducing a symmetric loss function with respect to positive and negative data, leading to improved convergence stability. SFFA (Terres-Escudero, Ser, and Bringas 2025) further enforces symmetry by dividing each layer’s neurons into positive and negative subsets, selectively activating them based on the input’s polarity. To overcome the lack of inter-layer interaction in FF, Layer Collaboration (Lorberbom et al. 2023) proposes a mechanism that allows goodness information to flow across layers, improving representational coherence and overall accuracy. However, these variants have not yet been extended to CNNs, limiting their applicability to more complex classification tasks.
@@
The spatial goodness vectors obtained as described above are used for layer-wise supervision via classification. A common approach is to employ auxiliary classifiers (Belilovsky, Eickenberg, and Oyallon 2019; Nøkland and Eidnes 2019); however, this introduces additional parameters and requires local backpropagation through the auxiliary branches, undermining the goal of achieving fully backpropagation-free training. To address this, we adopt a fixed random projection approach inspired by prior works (Huang, Zhu, and Siew 2006; Huang and Aminifar 2025; Apolinario, Roy, and Roy 2025). Specifically, each layer’s goodness vector is projected into the $N$-dimensional class space via a linear transformation defined by a weight matrix $\mathbf{W}_{l}\in\mathbb{R}^{\left(C_{l}\cdot P_{l}^{2}\right)\times N}$ and a bias vector $\mathbf{b}_{l}\in\mathbb{R}^{1\times N}$, both of which are randomly initialized at the beginning of training and remain fixed throughout. The generation procedure for these weight and bias parameters is as follows:
$\mathbf{W}_{l}\sim\mathcal{N}\left(0,\frac{1}{\sqrt{N}}\right),\mathbf{b}_{l}\sim\mathcal{N}\left(0,\frac{1}{\sqrt{N}}\right)$ (4)
The scaling factor $1/\sqrt{N}$ for $\mathbf{W}_{l}$ follows the Johnson-Lindenstrauss lemma, ensuring that the $\ell_{2}$ norm of the projected vector remains close to that of the original goodness vector (Johnson and Lindenstrauss 1984). The bias vector $\mathbf{b}_{l}$ is similarly scaled by $1/\sqrt{N}$, serving as a static offset that improves the logit distribution after projection. The classification logits $\mathbf{a}_{l}\in\mathbb{R}^{1\times N}$ are obtained from $\mathbf{g}_{l}$ via:
@@
As illustrated in Figure 2, during one forward pass, each layer is trained independently using spatial goodness-based supervision. Specifically, the convolutional output at each layer is first passed through a ReLU nonlinearity. The resulting activated feature maps are then used to compute the layer’s spatial goodness and the corresponding loss, while simultaneously being forwarded to the subsequent operations—pooling (if present) and layer normalization (Ba, Kiros, and Hinton 2016). Critically, each layer’s output is detached before propagation to downstream layers, preventing gradient flow across layers and maintaining strict layer-wise independence.
@@
We evaluate our approach on five benchmark datasets: MNIST (Lecun et al. 1998), FashionMNIST (F-MNIST) (Xiao, Rasul, and Vollgraf 2017), CIFAR-10, CIFAR-100 (Krizhevsky and Hinton 2009), and ImageNet (ILSVRC 2012). To assess scalability, we include ImageNet’s 1.28 million training and 50,000 validation images across 1000 classes. For MNIST and F-MNIST, the original training sets are split into 50,000 training and 10,000 validation samples, and models are trained for 100 epochs. For CIFAR-10 and CIFAR-100, we use 45,000 training and 5,000 validation samples, with training conducted over 400 epochs. For all datasets except ImageNet, we adopt the AdamW optimizer with a weight decay of 0.001, a batch size of 128 and an initial learning rate of 0.0002, which is decayed to 0.00001 using a cosine annealing scheduler. For ImageNet, we use the SGD optimizer with momentum 0.9 and weight decay 0.0001. The batch size is 512, and the learning rate is scheduled from 0.05 to 0.0005 via cosine annealing. Models are trained for 45 epochs. Training is performed on 2 NVIDIA A100 GPUs (80 GB) and 1 Intel(R) Xeon(R) Gold 6338 CPU. We report the mean accuracy and standard deviation averaged over three independent runs. In each run, the model checkpoint with the highest validation accuracy is selected for final evaluation on the test set. Additional details, including model architectures, data augmentation strategies and so on, are provided in the supplementary materials.
@@
Table 4: Test accuracy (%) comparison: ASGE vs. standard BP, non-$F$ and FF-based BP-free methods.

### v2

@@ Abstract
The Forward-Forward (FF) algorithm offers a promising alternative to backpropagation (BP). Despite advancements in recent FF-based extensions, which have enhanced the original algorithm and adapted it to convolutional neural networks (CNNs), they often suffer from limited representational capacity and poor scalability to large-scale datasets, primarily due to exploding channel dimensionality. In this work, we propose adaptive spatial goodness encoding (ASGE), a new FF-based training framework tailored for CNNs. ASGE leverages feature maps to compute spatially-aware goodness representations at each layer, enabling layer-wise supervision. Crucially, this approach decouples classification complexity from channel dimensionality, thereby addressing the issue of channel explosion and achieving competitive performance compared to other BP alternatives. ASGE outperforms all other FF-based approaches across multiple benchmarks, delivering test accuracies of 99.65% on MNIST, 93.41% on FashionMNIST, 90.62% on CIFAR-10, and 65.42% on CIFAR-100. Moreover, we present the first successful application of FF-based training to ImageNet, with Top-1 and Top-5 accuracies of 51.58% and 75.23%. Furthermore, we propose three prediction strategies to achieve flexible trade-offs among accuracy, parameters and memory usage, enabling deployment under diverse resource constraints.
Codehttps://github.com/KAI-Laboratory/ASGE
Introduction
@@
Artificial Neural Networks (ANNs) have become a cornerstone of modern machine learning, largely owing to the effectiveness of the backpropagation (BP) algorithm (Rumelhart et al. 1986). BP has driven the success of ANNs across a diverse range of applications (Goodfellow et al. 2014; He et al. 2015; LeCun et al. 2015; Kingma and Ba 2017; Vaswani et al. 2023) by enabling efficient training through gradient-based optimization. The algorithm proceeds in two phases: a forward pass computes the network’s predictions and their associated error relative to the ground truth, followed by a backward pass that propagates this error through the network. This yields gradients for each layer’s parameters (weights and biases), which are then used by optimization methods such as stochastic gradient descent (SGD) (Bottou 2010) to iteratively minimize the prediction error. However, BP suffers from inherent limitations. It requires symmetric weight matrices for the forward and backward passes to compute parameter gradients via the chain rule (Grossberg 1987). Additionally, it necessitates the storage of intermediate activations until parameter updates occur (Czarnecki et al. 2017). These constraints incur substantial computational and memory overhead, while simultaneously restricting the degree of parallelization, ultimately rendering BP inefficient and ill-suited for deployment on resource-constrained hardware.
@@
In response, numerous backpropagation-free (BP-free) strategies have emerged to address these limitations, including Target Propagation (TP) (Bengio 2014), Feedback Alignment (FA) (Lillicrap et al. 2016), Direct Feedback Alignment (DFA) (Nøkland 2016), Deep Random Target Projection (DRTP) (Frenkel et al. 2021), and Recursive Local Representation Alignment (rLRA) (Ororbia et al. 2023), alongside other biologically inspired or local learning methods (Hebb 2002; Moraitis et al. 2022; Dellaferrera and Kreiman 2023). Although these approaches alleviate some of the BP limitations, they continue to face challenges in balancing the trade-offs between prediction accuracy, memory efficiency and computational cost.
@@
A recent promising alternative is the Forward-Forward algorithm (FF) (Hinton 2022), which employs a layer-wise training strategy through dual forward passes: one with positive samples and the other with negative samples. Each layer independently updates its parameters to maximize a “goodness” metric for positive inputs while minimizing it for negative ones. By eliminating the need for backward passes, FF significantly reduces memory overhead and facilitates pipe-lined parallelism. However, Figure 1 illustrates a critical limitation of FF when applied to convolutional neural networks (CNNs): prediction accuracy tends to deteriorate progressively with network depth. This indicates a weakening capacity for feature abstraction in deeper layers, thereby constraining the algorithm’s effectiveness in deep architectures and on complex tasks. Recent enhancements, such as Channel-wise Competitive (CwC) learning (Papachristodoulou et al. 2024) and DeeperForward (Sun et al. 2025), attempt to mitigate this issue. While partially effective, they introduce an undesirable coupling between channel dimensionality and classification complexity. Specifically, they require layer-wise channels $C_{l}$ to scale linearly with classes $N$, i.e., $C_{l}\in\Theta(N)$. This scaling becomes impractical for large-scale datasets like ImageNet (Russakovsky et al. 2015) due to the dramatic increase in channel dimensionality.
@@
BP (Rumelhart et al. 1986) enables end-to-end optimization by propagating errors from the output layer back through the network. However, it suffers from several inherent limitations. First, it requires weight symmetry between the forward and backward passes (Grossberg 1987). Second, BP introduces the update locking problem (Czarnecki et al. 2017), wherein weight updates cannot be performed until the entire forward pass is completed. These constraints result in high memory overhead due to the need to store intermediate activations and hinder layer-wise parallelism, ultimately rendering BP inefficient for deployment on resource-constrained hardware.
@@
To overcome the limitations of backpropagation, numerous alternatives have been proposed that either avoid or approximate its core mechanism. Feedback Alignment (FA) (Lillicrap et al. 2016) eliminates the requirement for weight symmetry in the backward pass by using fixed, randomly initialized feedback weights. DFA (Nøkland 2016) extends this by propagating output layer errors to hidden layers, bypassing the backward pathway altogether. Another related method, rLRA (Ororbia et al. 2023) enables weight updates without computing global gradients by recursively aligning each layer’s output with a synthetic target derived from the layer above. PEPITA (Dellaferrera and Kreiman 2023) replaces the backward pass with a second forward pass, in which the input is modulated according to the output-layer errors from the first forward pass. Notably, these approaches still necessitate full forward passes to compute global errors or store intermediate activations prior to updating weights. To eliminate update locking, local learning rules—which allow layers to update independently without waiting for the entire forward pass—are desirable. DRTP (Frenkel et al. 2021) achieves this by directly projecting true labels to each layer, generating layer-wise error signals that support local updates without waiting for the full forward pass. Another approach is Hebbian learning (Hebb 2002), a biologically inspired local learning rule that strengthens synaptic connections when pre-synaptic and post-synaptic neurons are activated simultaneously. Despite its biological plausibility, Hebbian learning does not support supervised learning or error correction. SoftHebb (Moraitis et al. 2022; Journé et al. 2023) extends Hebbian learning by incorporating soft winner-take-all dynamics and differentiable surrogate gradients, enabling the training of deeper networks while preserving the characteristics of local learning. Similarly, Predictive Coding (Rao and Ballard 1999; Millidge et al. 2022) offers a local learning framework in which each layer predicts the activity of the layer below, using the resulting prediction error to locally update both prediction and representation neurons.
@@
The FF algorithm (Hinton 2022) introduces a fundamentally different training paradigm that entirely avoids the need for backpropagation. It employs a layer-wise learning mechanism in which each layer is independently supervised through two distinct forward passes—one with positive data and the other with negative data. Each layer computes a “goodness” score, defined as the sum of squared activations, and is trained using a binary cross-entropy loss to classify inputs as positive if the goodness exceeds a predefined threshold, and negative otherwise. Several extensions have been proposed to enhance the performance of FF. Contrastive Forward-Forward (Aghagolzadeh and Ezoji 2025) redefines goodness as the dot-product similarity between representations of different samples and introduces layer-wise supervised contrastive learning to improve the discriminative power of the network. SymBa (Lee and Song 2023) addresses the inherent asymmetry in binary cross-entropy loss by introducing a symmetric loss function with respect to positive and negative data, leading to improved convergence stability. SFFA (Terres-Escudero et al. 2025) further enforces symmetry by dividing each layer’s neurons into positive and negative subsets, selectively activating them based on the input’s polarity. To overcome the lack of inter-layer interaction in FF, Layer Collaboration (Lorberbom et al. 2023) proposes a mechanism that allows goodness information to flow across layers, improving representational coherence and overall accuracy. However, these variants have not yet been extended to CNNs, limiting their applicability to more complex classification tasks.
@@
The spatial goodness vectors obtained as described above are used for layer-wise supervision via classification. A common approach is to employ auxiliary classifiers (Belilovsky et al. 2019; Nøkland and Eidnes 2019); however, this introduces additional parameters and requires local backpropagation through the auxiliary branches, undermining the goal of achieving fully backpropagation-free training. To address this, we adopt a fixed random projection approach inspired by prior works (Huang et al. 2006; Huang and Aminifar 2025; Apolinario et al. 2025). Specifically, each layer’s goodness vector is projected into the $N$-dimensional class space via a linear transformation defined by a weight matrix $\mathbf{W}_{l}\in\mathbb{R}^{\left(C_{l}\cdot P_{l}^{2}\right)\times N}$ and a bias vector $\mathbf{b}_{l}\in\mathbb{R}^{1\times N}$, both of which are randomly initialized at the beginning of training and remain fixed throughout. The generation procedure for these weight and bias parameters is as follows:
$\mathbf{W}_{l}\sim\mathcal{N}\left(0,\frac{1}{N}\right),\mathbf{b}_{l}\sim\mathcal{N}\left(0,\frac{1}{N}\right)$ (4)
The scaling factor $1/N$ for $\mathbf{W}_{l}$ follows the Johnson-Lindenstrauss lemma, ensuring that the $\ell_{2}$ norm of the projected vector remains close to that of the original goodness vector (Johnson and Lindenstrauss 1984). The bias vector $\mathbf{b}_{l}$ is similarly scaled by $1/N$, serving as a static offset that improves the logit distribution after projection. The classification logits $\mathbf{a}_{l}\in\mathbb{R}^{1\times N}$ are obtained from $\mathbf{g}_{l}$ via:
@@
As illustrated in Figure 2, during one forward pass, each layer is trained independently using spatial goodness-based supervision. Specifically, the convolutional output at each layer is first passed through a ReLU nonlinearity. The resulting activated feature maps are then used to compute the layer’s spatial goodness and the corresponding loss, while simultaneously being forwarded to the subsequent operations—pooling (if present) and layer normalization (Ba et al. 2016). Critically, each layer’s output is detached before propagation to downstream layers, preventing gradient flow across layers and maintaining strict layer-wise independence.
@@
We evaluate our approach on five benchmark datasets: MNIST (Lecun et al. 1998), FashionMNIST (F-MNIST) (Xiao et al. 2017), CIFAR-10, CIFAR-100 (Krizhevsky and Hinton 2009), and ImageNet (ILSVRC 2012). To assess scalability, we include ImageNet’s 1.28 million training and 50,000 validation images across 1000 classes. For MNIST and F-MNIST, the original training sets are split into 50,000 training and 10,000 validation samples, and models are trained for 100 epochs. For CIFAR-10 and CIFAR-100, we use 45,000 training and 5,000 validation samples, with training conducted over 400 epochs. For all datasets except ImageNet, we adopt the AdamW optimizer with a weight decay of 0.001, a batch size of 128 and an initial learning rate of 0.0002, which is decayed to 0.00001 using a cosine annealing scheduler. For ImageNet, we use the SGD optimizer with momentum 0.9 and weight decay 0.0001. The batch size is 512, and the learning rate is scheduled from 0.05 to 0.0005 via cosine annealing. Models are trained for 45 epochs. Training is performed on 2 NVIDIA A100 GPUs (80 GB) and 1 Intel(R) Xeon(R) Gold 6338 CPU. We report the mean accuracy and standard deviation averaged over three independent runs. In each run, the model checkpoint with the highest validation accuracy is selected for final evaluation on the test set. Additional details, including model architectures, data augmentation strategies and so on, are provided in the supplementary materials.
@@
ASGE(ours) ResNet18-CHx4 - - - - 51.58 75.23
[…]